OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5 (−10, −5) (−7, −2) (7, 2) (5, 10)

OpenStudy (anonymous):

@dumbcow

OpenStudy (whpalmer4):

okay, you can solve this by substitution. y^2 + x^2 = 53 y - x = 5 solve the second equation for x or y. plug the result into the first equation. that should give you an equation in just one variable. solve it.

OpenStudy (anonymous):

Two negatives make a positive

OpenStudy (whpalmer4):

@MountainDew is guessing. you need to do what I said.

OpenStudy (anonymous):

I think its a because -10 - -5= 5

OpenStudy (anonymous):

Im not guessing I had to look for a second

OpenStudy (anonymous):

i tried to understand, and do the equations but i came out with a fraction

OpenStudy (whpalmer4):

here's how you can check your answer. suppose it is (-10, -5). that means both equations must be correct if you substitute x = -10 and y = -5. Are they? y - x = 5 (-5) - (-10) = 5 -5 + 10 = 5 well, okay, so far, so good y^2 + x^2 = 53 (-5)^2 + (-10)^2 = 53 25 + 100 = 53 uh, oops.

OpenStudy (whpalmer4):

A is NOT correct.

OpenStudy (anonymous):

i was thinking it was (5,10)

OpenStudy (anonymous):

but it came out the same as -10, -5

OpenStudy (anonymous):

no more help?

OpenStudy (whpalmer4):

here's how you solve it. y - x = 5 y^2 + x^2 = 53 solve the first one for y: y - x= 5 y = x + 5 substitute that into the other equation: y^2 + x^2 = 53 (x+5)^2 + x^2 = 53 x^2 + 5x + 5x + 25 + x^2 = 53 2x^2 + 10x + 25 = 53 2x^2 + 10x - 28 = 0 x^2 + 5x - 14 = 0 factoring (x+7)(x-2) = 0 solutions are x = -7, x = 2 so at x = -7, y = x+5 = -2 (-7,-2) is a solution to the system of equations

OpenStudy (anonymous):

7.2 is the answer

OpenStudy (anonymous):

i then tried 7,2 AND THE first equation did not equal out to be 5

OpenStudy (anonymous):

7-2

OpenStudy (anonymous):

oh the negative 2 ok i see now

OpenStudy (whpalmer4):

another solution is x = 2, y = x+5 = 7 or (2, 7) again checking: x=2, y = 7 x^2 + y^2 = 53 2*2 + 7*7 = 53 4 + 49 = 53 53 = 53

OpenStudy (anonymous):

=5

OpenStudy (jagatuba):

@whpalmer4 gave you the correct answer and the proper method for solving.

OpenStudy (anonymous):

(−10, −5) (−7, −2) (7, 2) (5, 10) those are my possible answers

OpenStudy (anonymous):

I said it was b -_-

OpenStudy (whpalmer4):

it's a quadratic equation (has an x^2 in it), so there are two solutions. only one of them is in the answer list. @MountainDew you were guessing based on your "two negatives make a positive". that would also apply to A, and that was not a solution.

OpenStudy (anonymous):

So??

OpenStudy (anonymous):

Sorry MountainDew, they say youre wrong :(

OpenStudy (anonymous):

I was just trying to help I said maybe

OpenStudy (whpalmer4):

so, the idea here is to show @eliteent how to solve these problems, not how to guess at the answer and get it wrong. doing it correctly is not difficult.

OpenStudy (anonymous):

Well at least I tried to help

OpenStudy (anonymous):

bro you did good, dont worry about it.

OpenStudy (whpalmer4):

no one is disputing that you tried to help.

OpenStudy (anonymous):

Okay bye

OpenStudy (jagatuba):

This was actually a good question thread because it wasn't just he asker that learned something. :)

OpenStudy (anonymous):

Yeah, very nice this website is so elaborately set up. Its very helpful

OpenStudy (anonymous):

(-7, -2), (2, 7)